Fermi National Laboratory


Accelerator Update for October 3, 2005 - October 5, 2005

Monday October 3
The day shift began with Operations monitoring store 4428 and a Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with NuMI off due to a Horn ground fault, and with the MiniBooNE and the Meson MCenter experiments taking beam.

Operations began preparing the Recycler at 9:30 AM to stash antiprotons. The stash was completed by 11:14 AM. Pbar resumed stacking and MCenter resumed taking beam.

At 3:51 PM, the Recycler suffered a vacuum burst and had an emittance blowup.

The Recycler began preparing to stash around 4 PM. The stash ended and Pbar resumed stacking at 6:04 PM. MCenter resumed taking beam.

Operations terminated store 4428 at 11:03 PM.

Operations prepared CDF for a collision hall access at 11:17 PM.

Cryo system technicians began a Tevatron (TeV) sector D1 wet engine repair at 11:28 PM.

Tuesday October 4
The midnight shift began with Operations monitoring the stash, with Pbar stacking, with NuMI off due to a Horn ground fault, and with MiniBooNE and MCenter taking beam.

A cryo tech reported at 1:42 AM that the D1 wet engine work was completed.

The Recycler began preparing to stash at 1:52 AM.

CDF came out of access at 2:12 AM.

The stash ended and Pbar resumed stacking at 2:44 AM.

Operations reported at 2:59 AM that the TeV was locked in a 90 GeV null ramp.

At 3:18 AM, Operations reported that they had run the TeV recovery program and a QBS ramp, but the TeV was still stuck at a 90 GeV null ramp. Operators continued their investigations.

At 5:05 AM, Operations reported that they found a problem with the C4 UPS (Uninterruptible Power Supply) and then later a QPM problem. Operations contacted an expert and then changed out the C4 UPS.

Operations reported at 5:47 AM that a TeV expert had managed to get the QPM working. Operators began the TeV turn on.

Operations put the TeV through a dry squeeze at 6:26 AM.

Operations put the TeV into shot setup at 6:51 AM.

The Recycler began preparing to stash at 9:02 AM.

Operations established combination store 4431 at 9:11 AM with an initial luminosity of 141.28E30. This store not only set a New TeV Record Luminosity, but it also broke a World Record Luminosity (for a hadron collider) of ~140E30 set at CERN's ISR in 1982.

Operations prepared Pbar for a tunnel access at 10:22 AM.

Operations prepared Linac and the I- Source for access at 10:26 AM.

Experts came out of the I- Source access at 10:39 AM.

Operations reported that the Linac access had ended and that Operators had sent beam to the Linac dump by 1 PM.

MiniBooNE resumed taking beam at 1:42 PM and Pbar resumed stacking at 1:49 PM.

MCenter resumed taking beam at 2:14 PM.

Recycler began preparing to stash around 8:30 PM. The stash ended and Pbar resumed stacking at 11:18 PM. MCenter also resumed taking beam.

Wednesday October 5
The midnight shift began with Operations monitoring store 4431 and the stash, with Pbar stacking, with NuMI off due to a Horn ground fault, and with MiniBooNE and MCenter taking beam.

Recycler began preparing to stash at 3:53 AM. The stash ended and Pbar resumed stacking at 5:07 AM.

At 6:08 AM, the Booster East Anode Power Supply (EAPS) and the BRF4 tripped off. Operations restored beam by 6:19 AM, but BRF4 remained off. Operations brought BRF4 back on line by 6:50 AM.

The Plans for Wednesday
The plans for today are to stack, stash, and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/5/2005   email Fermilab